OpenStudy (anonymous):

Evaluate: 52 ÷ (4 + 1) 1 3 4 5

OpenStudy (anonymous):

52/(5) = ? It's not any of the answers above

OpenStudy (anonymous):

well thats what it says on my homework

OpenStudy (anonymous):

Is it 52/(4x+1) ?

OpenStudy (anonymous):

no its exacly like that

OpenStudy (anonymous):

52/5 = 10.4

OpenStudy (anonymous):

thats wrong

OpenStudy (anonymous):

5(10.4)=52.

OpenStudy (anonymous):

nope ill just give you another one

OpenStudy (anonymous):

Maybe it's a typo on the homework

OpenStudy (anonymous):

Evaluate the following expression using the values given: Find 5x - 3y - z if x = -2, y = 2, and z = -3 -19 19 13 -13

OpenStudy (anonymous):

i think idk

OpenStudy (anonymous):

replace x with -2, y with 2, and z with -3

OpenStudy (anonymous):

5(-2)-3(2)-(-3)

OpenStudy (anonymous):

its 13

OpenStudy (anonymous):

-13.

OpenStudy (anonymous):

so close

OpenStudy (anonymous):

Evaluate: -3 - (-4) -7 -1 1 -12

OpenStudy (anonymous):

-12